### v1

# Culturally Aware and Adapted NLP: A Taxonomy and a Survey of the State of the Art

Abstract
The surge of interest in culturally aware and adapted Natural Language Processing (NLP) has inspired much recent research. However, the lack of common understanding of the concept of “culture” has made it difficult to evaluate progress in this emerging area. Drawing on prior research in NLP and related fields, we propose an extensive taxonomy of elements of culture that can provide a systematic framework for analyzing and understanding research progress. Using the taxonomy, we survey existing resources and models for culturally aware and adapted NLP, providing an overview of the state of the art and the research gaps that still need to be filled.
1 Introduction
Culture is rapidly becoming an important research topic in Natural Language Processing (NLP), with a significant recent surge in the number of published papers (Figure 1). Given the keen interest in this area and its importance for the safety and fairness of Large Language Models (LLMs), it is now important to consolidate existing research on culturally aware and adapted NLP to take stock of the progress made so far and to identify research gaps. However, this is challenged by the fact that there is no common understanding of the concept of “culture” in NLP.
Prior work such as Hershcovich et al. (2022); Hovy and Yang (2021) laid the vital foundations for understanding ways in which language, culture and society interact in NLP. Hershcovich et al. (2022) proposed a simple taxonomy derived from the interaction between language and culture that captures some elements of culture (linguistic form and style, objectives and values, common ground, and aboutness). Recent papers have adopted a diversity of other definitions (using e.g., “proxies of culture”,  Adilazuarda et al. 2024), making it challenging to assess progress in this area.
Figure 1: Papers published on arXiv (cs.CL) that contain the keyword “culture”, aggregated every 3 months.
Figure 2: An overview of the taxonomy.
Although culture is arguably a complex concept, we can establish a common understanding by drawing on the definitions developed within the humanities and social sciences. In these fields Tylor (1871); Kroeber and Kluckhohn (1952); White (1959); UNESCO (1982); Matsumoto and Juang (1996); Blake (2000), most definitions of culture encompass people, groups of people and the interactions among both individuals and groups. [^1] Differences between cultures can exist within groups, such as due to languages; however, variations in language do not imply cultural differences in certain aspects and vice-versa (e.g., United Kingdom versus United States).
In anthropology, White (1959) provides valuable insights into the locus of culture, identifying three key dimensions: 1) “within human” (such as concepts, traditions, beliefs, social practices, etc.), 2) between “social interaction among human beings,” and 3) outside of human but “within the patterns of social interaction” (in materialized objects such as tools, arts). An examination of such work reveals that not only people, but also social interaction are critical components of culture.
In NLP, understanding of culture could benefit from the granularity needed to distinguish between more nuanced differences among cultural elements. Additionally, although social interactions are important, they have not been explicitly defined or categorized in prior work on cross-cultural NLP Hershcovich et al. (2022).
We introduce a new extensive taxonomy of the elements of culture that addresses these problems. Expanding on the basic categories of prior work Hershcovich et al. (2022); Hovy and Yang (2021) and grounding the definition of culture in the relevant anthropology literature Tylor (1871); Kroeber and Kluckhohn (1952); White (1959), we introduce a fine-grained taxonomy that also covers social interactions. We then use this taxonomy to organize and analyze existing works in culturally aware and adapted NLP according to the elements of culture. Our survey of over a hundred publications in NLP provides an up-to-date view of cultural adaptation resources and models and identifies areas of progress as well as research gaps. We hope our taxonomy and analysis can enable and inspire future research in this emerging area.
2 The Taxonomy
In this section, we present our new taxonomy of culture. It differs from earlier NLP works Hershcovich et al. (2022); Hovy and Yang (2021) that aimed to define elements of culture in that (i) it is grounded in well-established elements of culture in anthropology literature Tylor (1871); Kroeber and Kluckhohn (1952); White (1959), (ii) consists of more fine-grained elements than in earlier work, and (iii) allows for a wider consideration of how social factors and variations in humans influence culture.
The taxonomy – shown in Figure 2 has two main branches: cultural elements3.1) within humans and outside of humans but within the social context White (1959) and sociocultural elements3.2) which affect and are affected by cultural elements through social interactions and communication among humans. [^2] Note that all elements of culture may interact and/or affect each other depending on context. Our taxonomy abstracts away from this contextual variation.
Cultural Elements. Cultural elements are based on well-established definitions of culture Tylor (1871); Kroeber and Kluckhohn (1952); White (1959):
Concepts: basic units of meaning underlying objects, ideas, or beliefs Jackendoff (2012).
Knowledge: information that can be acquired through education or practical experience.
Values: beliefs, desirable end states or behaviours ranked by relative importance that can guide evaluations of things Schwartz (1992).
Norms and Morals: set of rules or principles that govern people’s behaviour and everyday reasoning Cialdini et al. (1991); Hechter and Opp (2001); Bicchieri et al. (2018); Gert and Gert (2002).
Linguistic Form: “how” to construct an utterance (Hershcovich et al., 2022, e.g., dialects).
Artifacts: “materialized” items as the productions of human culture, they can be forms of art, tools, machines White (1959) etc.
Sociocultural Elements. “Sociocultural” elements refer to social factors within the scope of NLP that influence and are influenced by previously discussed elements. Leveraging the work of Hovy and Yang (2021), we identify relevant elements:
Relationship: connection between two or more individuals or groups (e.g., father-son, colleagues).
Context: the “containers” of communications Yang (2019), which can be linguistic such as surrounding sentences or extra-linguistic Hovy and Yang (2021) including social settings (e.g., at a wedding), non-verbal cues (e.g., gesture), or historical contexts (e.g., colonization).
Communicative Goals: the intention behind language use (e.g., requests, apologies).
Demographics: the characteristics of people, such as economic income, education level, nationality, location, political view, family status, etc.
3 Elements of Culture in Current NLP (Resources) Literature
In this section, we survey and categorize NLP resources published in leading *CL venues since 2022 (details in Appendix A) according to our new taxonomy (§2). For an overview of papers published in each cultural element class of the taxonomy, see Table 1 in the Appendix.
We observe that in resources, culture can be captured in 1) the data itself, or 2) in the labels (e.g., multi-culturally annotated).
3.1 Cultural Elements
3.1.1 Concepts
We can divide concepts into 1) basic concepts that are “configured” differently, reflecting the cultural-specific way of thinking, [^3] For example, one can explore the citizen science project for lexicon associations: https://smallworldofwords.org/en/project/home. and 2) concepts that are unique to a culture Wierzbicka (1992). [^4] For example, “Kopi Ga Dai” in Singaporean English versus “double-double” in Canada, both referring to coffee with extra sweetness and creaminess, but very different.
Recent NLP works have studied the grounding of time expressions to specific hours across different cultures, as detailed in Shwartz (2022, TimeExpressions, 27 languages). Cao et al. (2024b, CulturalRecipes, 2 languages) study culinary and food-related concepts within the context of recipe adaptations. Majewska et al. (2023); Hu et al. (2023) (4 languages each) culturally adapt names and locations for multilingual Task-Oriented Dialogue (TOD) datasets to enable cultural adaptation of TOD systems. Nevertheless, both training and evaluation datasets are still lacking in this area for diverse cultures, languages, and concept categories (e.g., rituals, aesthetics, kinship, spatial relations).
NLP research has also investigated the usage of concepts spanning multiple categories across cultures, e.g., through metaphors and other figurative expressions Kabra et al. (2023) or traditional proverbs and sayings Liu et al. (2023b).
In vision and language (VL) settings, culturally significant or unique concepts have been incorporated into reasoning and captioning tasks such as in MaRVL Liu et al. (2021), GD-VCR Yin et al. (2021), and XM3600 Thapliyal et al. (2022). Additionally, these concepts have been employed to evaluate multimodal content adaptations Khanuja et al. (2024) or to generate multicultural images Liu et al. (2023e); Ventura et al. (2023) with text-to-image models. These datasets are limited in size (e.g., due to the high cost of annotations) and most are available for evaluation only.
3.1.2 Knowledge
Cultural knowledge can be factual or common sense. [^5] Common sense and norms are sometimes used interchangeably in NLP. Norms are acceptable behavioural patterns of a group (§2), which we will discuss in §3.1.4. What weather phenomena can be expected if a rapidly rotating tropical storm forms off the coast of our country? (It’s likely called a hurricane in the US and a typhoon in Korea.) Is tofu pudding sweet or savoury by default? (In China, tofu pudding is typically sweet in the south and savoury in the north.)
We identified three major types of resources in NLP literature: 1) probing (by masking entities), 2) multiple choice question answering (MCQA), and 3) knowledge bases.
mLAMA Kassner et al. (2021), GeoMLAMA Yin et al. (2022), DLAMA Keleg and Magdy (2023), and FMLAMA Zhou et al. (2024) probe the diversity of knowledge in LMs.
Recently, MMLU Hendrycks et al. (2021) style QA benchmarks have greatly aided LLM development, inspiring many cultural variants of the MMLU suite. These suites include ArabicMMLU Koto et al. (2024a), CMMLU Li et al. (2023b), IndoMMLU Koto et al. (2023), SeaEval Wang et al. (2023a), KMMLU Son et al. (2024), CLIcK Kim et al. (2024), just to name a few. The cultural variants of MMLU typically cover some aspects of food, history, literature, geographical knowledge of countries, etc. in respective languages. However, MMLU-style benchmarks are typically constructed using exams and textbooks and cannot combine with other elements in culture. In contrast, common sense datasets (COPAL-ID, Wibowo et al. 2023 or IndoCulture, Koto et al. 2024b) can consider dialects and/or geographical diversity.
Lastly, CANDLE Nguyen et al. (2023), CultureAtlas Fung et al. (2024), MANGO Nguyen et al. (2024) are recent cultural knowledge bases created from either Wikipedia articles or using model distillation. They can be integrated with models to enhance models’ cultural awareness. However, diverse sources (i.e., other than Wikipedia, such as from local encyclopedia) should be considered to improve the coverage of knowledge bases.
3.1.3 Values
Diverse ranking of values among groups can result in differences in aboutness, communication styles, perceptions and multiple other dimensions Hofstede (1984, 2011). Such differences in pre-training data can be reflected in LLMs.
Many recent studies on evaluation (Johnson et al., 2022; Ramezani and Xu, 2023; Cao et al., 2023; Wang et al., 2023b; Durmus et al., 2023; Santurkar et al., 2023; Masoud et al., 2023; Havaldar et al., 2023, inter alia) show that LLMs align better with values of WEIRD (Western, Educated, Industrialized, Rich and Democratic, Henrich et al. 2010) people, raising concerns about the fairness and safety of LLM for others. Here, Pew Global Attitudes Survey (PEW) [^6] https://www.pewresearch.org/, the World Values Survey (WVS) [^7] https://www.worldvaluessurvey.org/ and the Hofstede Cultural Dimensions Hofstede (1984, 2011) are commonly used for evaluation, along with regional variants like the European Values Survey (EVS, EVS 2011). However, the questions of how to improve the model’s value alignment with diverse cultures, what resources to collect and whom to collect from remain unsolved Kirk et al. (2024).
Biases. In contrast to cultural values, biases have been long-studied in NLP, such as gender bias in machine translation (Stanovsky et al., 2019; Savoldi et al., 2021; Campolungo et al., 2022; Sandoval et al., 2023; Attanasio et al., 2023, inter alia) or bias towards particular social groups. Cultures can have different biases toward the same target groups or have unique biases that are more pronounced in certain cultures (e.g., castes, unnatural beauty standards).
To enable evaluations of cross-cultural variations in biases and develop transferable de-biasing methods, recent work has created culturally aware targets and attribute word sets. WEATHub Mukherjee et al. (2023) and CA-WEAT España-Bonet and Barrón-Cedeño (2022) are multilingual culturally aware extensions based on the standard WEAT Caliskan et al. (2017) categories. Alternatively, there are other cultural variants such as BIBED (identity-bias in Bengali, Das et al. 2023), CAMeL (Arabic and Western cultural entities, Naous et al. 2023), SeeGULL (regional and national identity-bias, Bhutani et al. 2024), SPICE (stereotypes in India in English, Dev et al. 2023), CHBias (gender/orientation/age/appearance biases, Chinese, Zhao et al. 2023) among others.
In addition to using word lists, bias evaluation datasets may also consist of sentence pairs (such as French CrowS-Pair, Névéol et al. 2022; WinoQueer, Felkner et al. 2023), biased sentence and context (KosBi, Lee et al. 2023a), conversational (CDIAL-BIAS, Zhou et al. 2022) or question answering (FORK: food-related customs, Palta and Rudinger 2023; SODAPOP, An et al. 2023; KoBBQ, Jin et al. 2024).
In general, there is considerable progress in cultural biases compared to other sub-areas. Many recent surveys on general biases are also available covering critical areas such as evaluations or de-biasing methods Meade et al. (2022); Dev et al. (2022); Sun et al. (2019); Delobelle et al. (2022), we refer readers to them for further details.
Hate. The perception of hatefulness in text varies across cultures, languages and dialects, as recently shown for hate speech classification by Zhou et al. (2023a, b); Lee et al. (2023c); Lwowski et al. (2022); Arango Monnar et al. (2022), among others. Such model disparities may be due to the labelling process, where annotations from diverse cultural groups could serve as a mitigation strategy. CREHate Lee et al. (2023b) provides hate speech annotations from various English-speaking countries, all applied to the same English dataset. This example is unique in recent research as it explores variations in perceptions within the same language, highlighting the need for further research.
Other Perceptions. The perception of politeness, aesthetic appeal or emotions can also vary across cultures House and Kasper (1981); Ringel et al. (2019); Masuda et al. (2008); Mesquita et al. (1997). For example, whether a piece of text is deemed humorous or ironic is culturally dependent. Frenda et al. (2023) tries to address this with a cross-cultural annotated irony corpora (EPIC) by people from five English-speaking countries. Similarly, visual elements in arts can elicit different emotions in different cultural groups. ArtELingo Mohamed et al. (2022) introduces a benchmark with Chinese, Arabic and Spanish captions and emotion labels for artworks, aimed at evaluating models’ cultural-transfer performance. This research area is significantly limited.
3.1.4 Norms and Morals
In ethics, a distinction is made between descriptive and normative morality Gert and Gert (2002). In NLP, this distinction is often overlooked Vida et al. (2023) with a greater emphasis on the “end product”, which is the final set of rules or principles and their judgements.
Several norm banks are available, constructed via automatic, semi-automatic or manual norm discovery from conversations, online resources such as Wikipedia, Reddit, government-aided websites or using crowd annotators Forbes et al. (2020); Fung et al. (2023); Moghimifar et al. (2023); Shi et al. (2024); CH-Wang et al. (2023); Ziems et al. (2023a, 2022b); Rai et al. (2024); Dwivedi et al. (2023). Norm banks have also been automatically adapted to defensible norms in fine-grained situations Pyatkin et al. (2023); Rao et al. (2023) or culturally aware inference tasks CH-Wang et al. (2023); Huang and Yang (2023) for LLM evaluation and adaptation.
Norms and morals can support downstream applications via e.g., enabling the alignment of models. MoralDial Sun et al. (2023) enables inquisition for moral alignment of LLMs through conversations. PROSOCIALDIALOG Kim et al. (2022) encourages more socially-aligned responses to problematic content. SocialDial Zhan et al. (2023), NormDial Li et al. (2023c) and RENOVI Zhan et al. (2024) aim to develop conversational experiences that better align with the expectations of a culture.
Nearly all resources are in English, sometimes Chinese, with a focus on Western, and occasionally Chinese or Indian, cultures.
3.1.5 Linguistic Form
Variations in linguistic form can be expressed 1) in the language itself, the habitual use of language, such as dialects; or 2) in style and is tightly coupled with other elements of culture (e.g., social norms) as well as sociocultural factors (e.g., pragmatics, Thomas 1983; Blum-Kulka and Olshtain 1984; Blum-Kulka 1987; Wierzbicka 2003; House and Kasper 1981).
Dialects. A dialect is a variant of a language Haugen (1966) at the local regional level (e.g., Hessian German), national level (e.g., Tunisian Arabic) or by other factors (e.g., African American Vernacular English vs. Standard American English).
NLP has predominantly focused on standard language, but there has been recent interest in dialects. Many works have focused on dialect identification Salameh et al. (2018); Abdelali et al. (2021); Yusuf et al. (2022); Hämäläinen et al. (2021), but how to enable LLMs to serve dialectal communities remains an open question.
Recently, Le and Luu (2023); Paonessa et al. (2023) identified the disparity of translation results within varieties of Vietnamese and German. Other research shows that models overestimate their performance on GLUE tasks Ziems et al. (2022a) and masked span predictions when tested with African American (Vernacular) English (AAE/AAVE) Deas et al. (2023). These dialect disparities correlate with linguistic, economic and social factors Kantharuban et al. (2023).
The available dialect datasets are typically translations between dialects and standard languages or produced via dialect normalization (in text only or in audio and text). Recent works include DialectEval Aepli et al. (2023), SDS-200 Plüss et al. (2022), STT4SG-350 Plüss et al. (2023) and the ones used in Kuparinen et al. (2023). Only some dialect studies exist for traditional generation tasks such as summarization or standard benchmark tasks (e.g., GLUE). DivSumm Olabisi et al. (2022) is a dialect-diverse tweets dataset with human-written extractive and abstractive summaries. In Held et al. (2023), a synthetic AAE GLUE corpus is created using a rule-based framework Ziems et al. (2023b). Recently, DIALECTBENCH Faisal et al. (2024) provides a benchmark for over 281 dialect variants by aggregating over existing datasets.
Overall, research on German and English dialects is more advanced (marginally) than other dialect types.
3.1.6 Artifacts
NLP research on artifacts has focused on (monolingual or mono-cultural) artifacts in texts, e.g., fairy tales, fiction, poetry and songs Xu et al. (2022); Jiang et al. (2023); Thai et al. (2022); Yang et al. (2019); Haider et al. (2020); Chakrabarty et al. (2021); Ou et al. (2023); Li et al. (2023a), or in multimodal such as movies, humour and memes Hong et al. (2023); Sharma et al. (2020); Liu et al. (2022a); Hessel et al. (2023), to name a few. While “artifacts” is an independent cultural element, usage in adaptation typically involves tasks that align with one or more previously mentioned categories. For example, in ArtELingo (in §3.1.3), the input data is on art but cross-cultural measurement studies perceptions, which is a part of values (and Impressions, Kruk et al. 2023 is another example using art). Similarly, translations of literary novels need to account for concept differences such as names Jiang et al. (2023) across cultures. Work on integrating cross-cultural differences in modelling and data acquisition with artifacts is limited.
3.2 Sociocultural Elements
Although cultural variations are easily identifiable through sociocultural elements, many current work use languages or countries as the boundary for divisions. These elements are understudied in current culturally aware NLP.
3.2.1 Relationship
In many cultures, concepts and forms of communication differ depending on the relationship between the speakers. For example, Chinese has distinct terms for elder vs. younger siblings. Translations to (and from) a language without this property may result in a loss of nuances in meaning. In Korea and Japan, misused politeness level in conversation can violate cultural norms Matsumoto (1988); Ambady et al. (1996), especially in different social relationships. Considering relationships is important for building resources and modelling culturally appropriate methods. Zhan et al. (2023, 2024) serve as recent examples with this consideration.
3.2.2 Context
Hovy et al. (2020); Akinade et al. (2023); Stewart and Mihalcea (2024) show that machine translation systems can fail without appropriate consideration of linguistic context, revealing its importance in resource and model development. Extra-linguistic contexts (e.g., settings) have proved especially important in conversational tasks Zhan et al. (2023), norm bank Ziems et al. (2023a) construction, and VL applications.
3.2.3 Communicative Goals
Different cultures can have distinctive communication styles depending on communicative goals. For example, people may use indirect language for refusal (vs. direct refusal with a “no”) to avoid confrontation House (2005). Cultures may also exhibit variations in responses to the same situation (e.g., how to make requests and when to apologize, Blum-Kulka and Olshtain 1984). Taking this type of variation into account is important for cross-cultural pragmatic-inspired tasks – an area that remains understudied.
3.2.4 Demographics
A household with a monthly income of less than 50 US dollars is likely to have different household items than that with 5000 US dollars Rojas et al. (2022). Névéol et al. (2022) also found that the original English CrowS-Pair dataset relied on names as proxies for a sociodemographic group (“Amy for women, Tyrone for African American men”,  Névéol et al. 2022), whereas the French version features direct references to sociodemographic groups. These data differences may stem not only from cultural influences but also from the demographics of the data contributors. Where and from whom one collects data matters, as it can result in dramatic differences in data and modelling.
Demographic information is also important in annotation Sap et al. (2022); Pei and Jurgens (2023); Santy et al. (2023), where a piece of text can be humorous to some people but offensive to others Meaney (2020). In such cases, culture may exist in the labels rather than in the data. Recently, Lee et al. (2023b); Frenda et al. (2023) show how to capture different cultural views of annotators using the same dataset.
4 Culturally Aware Resource Acquisition
Resources like those discussed in §3 are critical for developing culturally aware NLP. As additional resources are much needed, this section surveys methods for creating new resources.
Resources can be classified based on their acquisition methods—manual, automatic, or semi-automatic—and their source types: 1) new ones (New), or 2) culturally adapted existing ones (CA, e.g., through translation from the original data, followed by culturally appropriate changes). 1) can capture unique cultural phenomena but this is not always possible due to limited funds or access to experts or native speakers. 2) can offer an alternative, but an accurate account of cultural phenomena may prove more challenging.
4.1 Manual: Incorporating Native Speakers, Communities, and Experts
A common strategy is to employ native speakers or experts (e.g., professional translators or students) for data acquisition. This can be done via crowd-sourcing platforms such as Amazon Mechanical Turk and Prolific Liu et al. (2021, 2023b) or in a community-driven manner, leveraging networks such as Masakhane  [^8] https://www.masakhane.io/, IndoNLP  [^9] https://indonlp.github.io/, university mailing lists, or Slack/Discord of organizations.
New: Most existing culture resources have been built by involving native speakers or communities for dataset acquisition. Examples include Liu et al. (2021); Koto et al. (2023); Kabra et al. (2023); Liu et al. (2023b), among others. For non-language related communities, WinoQueer Felkner et al. (2023) utilizes diverse channels (such as Slacks/Discord, gay Twitter) to reach the LGBTQ+ community and generates bias benchmarks based on community survey results.
CA: When starting from existing datasets, some works have involved communities (e.g., using surveys) to determine the needed modifications and supplements to datasets Névéol et al. (2022); Jin et al. (2024); Hu et al. (2023).
In general, native speakers are typically consulted throughout the life-cycle of new data acquisition (from annotations to quality checks). However, the entire community is rarely consulted during the initiation stage (i.e., designing tasks). Involving native speakers may be costly and difficult sometimes, but the quality they bring is invaluable.
4.2 Automatic: Models and Pipelines
Since manual adaptation is slow and hard to scale, the use of automation has gained popularity in resource acquisition.
New: For instance, CANDLE Nguyen et al. (2023) proposes a pipeline to extract cultural commonsense knowledge using various techniques like NER extraction, cultural facet classification, concepts extraction and ranking through algorithms or LMs. NormsSAGE Fung et al. (2023) utilizes LLMs for norm discovery from conversation data, then performs model self-verification to validate and filter the data. CultureAtlas Fung et al. (2024) extracts cultural knowledge from Wikipedia and hyperlinked document pages using LLMs for filtering and adversarial knowledge generation.
Recent works have also used sociodemographic prompting Deshpande et al. (2023); Beck et al. (2024); Hwang et al. (2023); Santurkar et al. (2023) — extending input prompts with sociodemographic information — to generate outputs tailored to specific groups. Further research could potentially lead to a reduction of effort in data acquisition in specific areas, such as creating subcultural variations of data in WEIRD people. However, it has also been argued that LLMs do not accurately mimic individual or group behaviours Argyle et al. (2023); Aher et al. (2023); Beck et al. (2024).
CA: Putri et al. (2024) investigate automatic adaptation of Commonsense QA in Indonesian and Sundanese, where the adaptation process includes paraphrasing and concepts/names replacements. Current GPT models show disparities in automatic cultural adaptations for different languages, indicating the need for further research.
4.3 Semi-Automatic: Structured Resources, Model-in-the-Loop
As demonstrated by Putri et al. (2024), LLMs struggle with fully automated cultural adaptations. Alternatively, semi-automatic approaches combine the quality of manual work with scalability.
New: Methods have been used to generate seed data for humans to clean and label, possibly iteratively. NormBank Ziems et al. (2023a) uses LLMs to generate an initial set of roles and behaviours for creating detailed norm candidates grounded in situations, to be annotated by humans. Similarly, CH-Wang et al. (2023) build a dataset of cross-cultural norms for Chinese using few-shot and Chain of Thought (CoT) prompting Wei et al. (2022) with GPT-3 Brown et al. (2020) to generate norm candidates, followed by human verification and editing. Liu et al. (2023b) utilize LLMs to generate conversations featuring proverbs in different languages. Since many proverbs are figurative, most conversations are revised by human native speakers. SeeGULL Bhutani et al. (2024) generates candidate multilingual associations using PaLM-2 Anil et al. (2023) and few-shot demonstrations, then performs human annotations.
CA: For dialects, Ziems et al. (2023b, Multi-Value) create a framework based on well-documented dialect differences in English by utilizing Electronic World Atlas of Varieties of English (Kortmann et al., 2020, eWAVE), which allowed adaptation and creation of dialectic datasets that span 50 English dialects. Multi-Value was used in adapting standard corpus to dialectal variants in Held et al. (2023); Xiao et al. (2023). However, such structured resources for adaptations may not exist or be suitable for other elements of culture (e.g., for concepts, consistently replacing ‘bread’ with ‘rice’ would not be desirable).
5 Creating Culturally Adapted Models
Most work in culturally aware NLP has so far focused on resource creation and evaluation. Work on creating culturally adapted models is still emerging. In this section, we survey the methods that have been proposed to date for creating culturally adapted models from pre-trained (L)LMs. Figure 3 shows an overview of the approaches adopted so far with example approaches. Some individual works may incorporate more than one adaptation method for comparison; we may discuss works in relation to the main method adopted.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2cm, s sep=4pt, calign=edge midpoint, l sep=7.5pt, , [, phantom [Model §5, finetuning_pt [Data Aug., finetuning [Li and Zhang (2023), finetuning_work] ] [Cont. Pre-training; Auxilary Losses, finetuning [Hofmann et al. (2024); Li et al. (2024), finetuning_work] ] [Information Integration, finetuning [Yao et al. (2023); Friedrich et al. (2023), finetuning_work] ] [Parameter-Efficient Adaptations, finetuning [Xiao et al. (2023); Liu et al. (2023d), finetuning_work] ] ] ]
Figure 3: Categorization of the adaptation modelling methods and examples in each category.
5.1 Data Augmentation
It can be challenging to acquire large corpora for supervised training in cultural adaptation. Data augmentation has been supporting this process and increasing the robustness of models. Li and Zhang (2023) present a data augmentation technique (CultureMixup) for multilingual multicultural VL reasoning tasks. CultureMixup creates code-mixed data with concepts in English replaced through cultural concept mapping for supervised training. The cultural concept sets (for concept mapping) are constructed by querying hyponyms, synonyms, and hypernyms in the ConceptNet Speer et al. (2017) and WordNet Miller (1995). However, the optimal resource will depend on the cultural element in question (§3.1). For example, a cultural knowledge base might be better for norms adaptations.
5.2 Continual Pre-training, Auxiliary Losses
Continual pre-training (CPT, including instruction tuning), intermediate task training, and utilizing multi-task or auxiliary losses have also been used for cultural adaptation. CPT involves tuning a pre-trained LM with an additional unlabeled domain or language corpus before downstream task fine-tuning. It can improve downstream task performance via full-parameter training Xu et al. (2019); Han and Eisenstein (2019); Gururangan et al. (2020) or by training a small set of additional parameters while keeping the original model frozen Pfeiffer et al. (2020); Wang et al. (2021); Ke et al. (2022).
Recently, Hofmann et al. (2024) show that when combined with a geo-location prediction loss, CPT can help to increase the awareness of dialectal variations of pre-trained LMs. Wang et al. (2024) show that mining instructions containing cultural knowledge and performing instruction tuning can improve models’ ability in cultural knowledge reasoning. In VL settings, Bhatia and Shwartz (2023) use a cultural common sense knowledge graph from  Nguyen et al. (2023) for CPT, developing a geo-diverse LM for downstream commonsense reasoning tasks. CultureLLM Li et al. (2024) reformulates the World Value Survey as an intermediate training task. It shows that LLMs can enhance their downstream performance in multicultural hate speech classification where diverse value perceptions across cultures are required. However, adapting LLMs to culture may result in catastrophic forgetting (McCloskey and Cohen 1989, or termed “alignment tax” due to RLHF tuning, Ouyang et al. 2022; Askell et al. 2021), potentially worsening their performance on general tasks. Similarly, general continual fine-tuning may shift existing cultural values Choenni et al. (2024) in pretrained models. These warrant further investigation.
5.3 Other Forms of Information Integration
Missing cultural context, knowledge or demographic information have also been integrated into models in other ways:
As representations. Cao et al. (2024a) introduce the cuDialog dataset and models for multi-turn dialogue classification and prediction through a cultural lens. Using an encoder-decoder Transformer (such as mT5,  Xue et al. 2021), the cultural dimensions vectors are concatenated to the hidden states at each layer. The cultural context vectors were obtained through a regression task over the encoder hidden states with dialogue inputs to predict the corresponding cultural scores (based on Hofstede Culture Dimensions, Hofstede 1984).
As part of the prompt. AlKhamissi et al. (2024) enhance LLMs’ cultural alignment by integrating anthropological reasoning with demographic-based prompts. In a game setting with multicultural players,  Shaikh et al. (2023) encode sociocultural attributes into the input to enrich the model’s ability for game plays. Similarly, Yao et al. (2023) propose strategies to enhance the cultural translation ability of LLMs with reference-explanation of cultural entities. LLMs effectively leverage indirect descriptions obtained from external data sources or generated by LLMs as a prior step.
Friedrich et al. (2023) propose a retrieval-based method for moral reasoning tasks. Moral contexts specific to a culture (or user) are stored in a retrieval engine. When querying with moral questions, relevant contexts are retrieved and prepend to the query as input to LLMs. LLMs can change their default answer based on the retrieved context. Using a retrieval-augmented system can be a promising future direction for adapting rich and evolving cultural information.
5.4 Parameter-Efficient Adaptations
As LMs grow larger, parameter-efficient fine-tuning methods (i.e., PEFT, by fine-tuning a small number of parameters, such as the bottle-neck adapters, Houlsby et al. 2019; LoRA, Hu et al. 2022 etc.) become increasingly important for task adaptations. Given their success in cross-lingual transfer learning (Pfeiffer et al., 2020; Ansell et al., 2021; Liu et al., 2023a, c; Üstün et al., 2020, among others), PEFT can be a natural choice for cultural adaptation of e.g., dialects.
Recently, HyperLoRA Xiao et al. (2023) uses the Hypernetwork (Ha et al., 2017, a neural network for generating parameters) to generate LoRA adapters based on dialectal features. DADA Liu et al. (2023d) proposes to train a pool of dialectal linguistic feature adapters and dynamically compose the adapters for dialectal tasks. Being task agnostic, PEFT methods could prove important for cultural adaptations beyond dialects.
5.5 Outlook: Reinforcement Learning from Human Feedback (RLHF), Direct Preference Optimization (DPO)
The success of LLMs has popularized RLHF Christiano et al. (2017); Bai et al. (2022); Ouyang et al. (2022) and DPO Rafailov et al. (2023); Ivison et al. (2023) methods. RLHF fine-tunes LMs with feedback by fitting a reward model with human preferences, and then training a reinforcement learning-based policy to maximize the learned reward. DPO avoids RL training by using a simpler supervised learning objective for an implicit reward model.
Recent work shows that RLHF can enhance the performance of multilingual instruction tuning for LLMs Lai et al. (2023), while DPO can improve the multilingual reasoning abilities She et al. (2024) of LLMs. The use of RLHF or DPO for multilingual multicultural adaptation is still limited, but these examples suggest that the direction could be promising.
6 Further Discussions
As we have seen, “culture” emerges as a highly active area of NLP research. Yet significant work remains to be done on both resources and methods for various elements of culture.
An area that requires attention is the overall process of researching culturally aware NLP. First, it is important to consider whether and how the target communities could best benefit from cultural technologies at this point in time. For example, many dialects are primarily oral, and speech-to-speech or speech-to-text translations may be preferable over text-based applications Blaschke et al. (2024). Or, if the culture in question does not have LLMs with basic conversational fluency yet, it may be more important to address this problem first, before embarking on improving pragmatic alignments of LLMs for the culture. Getting the process right requires consultation with target communities Bird (2020); Liu et al. (2022b); Mager et al. (2023). Ethical data collection practices are also critical and technology ownership must be considered. We refer the readers to work such as Bird (2020); Smith (2021); Cooper et al. (2024) for further details of this.
Another area to consider is the integration of insights and practices from fields beyond NLP. Incorporating cultural factors and performing cultural adaptations has a long history in application areas, such as for video games O’hagan and Mangiron (2013), movies Pettit (2009), online learning systems Blanchard et al. (2005), and clinical psychology intervention Bernal et al. (1995); Barrera Jr and Castro (2006). Existing practices should be used as a foundation for adapting NLP applications to meet the needs of diverse cultural contexts.
7 Summary and Future Research Directions
Culturally aware and adapted NLP has recently emerged as a highly active research area. Significant progress has been made in the development of resources for capturing various elements of culture, but the development of NLP methods is still in its infancy. We will now summarize the main research gaps identified in this survey with respect to the categories of our new taxonomy (§2 for details):
Resources. Currently, resources exist for all elements of culture, with considerable progress made on values3.1.3, particularly in biases) and knowledge3.1.2, particularly for MMLU-style cultural knowledge benchmarks). However, research is lacking in the following areas:
Cultural: Multilingual data resources covering a diverse set of concepts (e.g., aesthetics, spatial relation) in both unimodal and multimodal (§3.1.1) for generation tasks are needed. Moreover, most recent developments in norms & morals are predominantly in English and reflect a monocultural perspective, which indicates the need for further development of multilingual and multicultural resources. Additionally, there is a significant gap in datasets that focus on different types of value perceptions (such as emotion and irony, §3.1.3), stylistic variations (§3.1.5), and artifacts (§3.1.6) across various cultural groups both in different languages and within languages.
Sociocultural: Resources that explicitly consider sociocultural elements of culture (§3.2) are still lacking. For example, collecting speaker relationships in dialogue datasets or distinguishing age groups in social norms datasets. These are needed to address the intricate relationship between culture and people in NLP.
Training Data and “CultureGLUE”: Most existing resources focus on evaluation, but there is a pressing need for training data. In addition, there is no unified cultural equivalent of the GLUE benchmark that encompasses all elements of culture across a wide range of cultures. Creating a multicultural “CultureGLUE” may be challenging at the moment, a good first step is to focus on individual cultures with diverse tasks and element coverage.
Modelling. While modelling methods for culture are generally under-explored, continual pretraining (§5.2) and prompting (§5.3) have received marginally more attention than other approaches. Research areas needing further exploration include:
PEFT-based Transfer Learning: Exploration of PEFT-based transfer learning techniques beyond dialects is limited (§5.4). Given their success in other NLP areas, these techniques warrant further investigation into other elements of culture, such as values or norms & morals.
RLHF, DPO and Other LLMs Specialties: Leveraging the success of LLMs, RLHF and DPO as a new promising avenue for cultural adaptation (§5.5). It would also be interesting to explore other unique abilities of LLMs (such as tool-use, role-playing etc.) for enabling culturally aware NLP.
Evolving Culture: Culture (slowly) evolves Boyd and Richerson (1988); Whiten et al. (2011), yet there have been little discussions on how to model and adapt to evolving culture. One potential approach is the use of retrieval-augmented systems to integrate evolving information (§5.3), which ensures models’ relevance to cultural shifts over time.
Overall - “Surface” versus “Deep” culturally adapted NLP. Resnicow et al. (1999) devise cultural adaptations for public health research into surface and deep adaptations, where the former considers familiar languages and concepts to the target groups, and the latter considers social and historical factors that influence the behaviours of the target groups. In NLP, surface adaptations might include using the same language as a culture and recognizing explicit cultural differences (e.g., asking LLMs “what is the meaning of …”). In contrast, deep adaptations might enable a model to “behave” (e.g., make decisions, pragmatically comply etc.) like a member of a culture without explicit inquisition (see Figure 4 in the Appendix for an illustration).
Only a few current works focus on adapting the behavioural aspect of models (which is becoming increasingly important with LLMs), and there has been no work to date measuring the depth and progress of cultural adaptations or when a model is fully culturally aware and culturally competent. Further research could explore these areas.
8 Conclusions
This paper proposes a new extensive taxonomy of culture that expands on earlier works in NLP and is grounded in well-established anthropology literature. The taxonomy provides a systematic framework for understanding and tracking progress in the emerging area of culturally aware and adapted NLP. We survey existing resources and methods in this area according to the taxonomy classes, identifying areas of strength as well as areas where research remains to be done. Our paper summarizes the state of the art and provides ideas for future development in this exciting and important area of research.
9 Limitations
In this work, we may only discuss research with respect to a single element of culture. Research related to culture can be multifaceted, considering multiple elements of culture simultaneously. Additionally, our survey emphasizes resources and adaptation methods due to the scope, without extensive coverage of evaluation and probing techniques.
Our taxonomy provides a systematic framework for understanding and tracking progress in the emerging area of culturally aware and adapted NLP. However, our taxonomy is not without its limitations. Future research could refine the taxonomy in areas like values or communicative goals, create subcategories and enhancing our understanding of how elements of culture interact (e.g., how changing values impact social norms over time).
Acknowledgements
This work was funded by the German Federal Ministry of Education and Research (BMBF) under the promotional reference 13N15897 (MISRIK). This work has also been supported by the UK Research and Innovation (UKRI) Frontier Research Grant EP/Y031350/1 EQUATE (the UK government’s funding guarantee for ERC Advanced Grants) awarded to Anna Korhonen at the University of Cambridge. Chen Cecilia Liu is supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
We thank Anjali Kantharuban, Shun Shao, and Anne Lauscher for an early discussion of different aspects of this work. We thank Ji-Ung Lee for feedback on a draft of this paper.
Appendix A Method
We prioritize main and findings papers from ACL, EMNLP, EACL, NAACL and TACL published in 2022, 2023 and early 2024. The search terms used are: “culture”, “cultural”, “geo-diverse”, “socio”, “social”, “moral”, “bias”, “norms” and we prioritize works that focus on datasets and modelling methods for cultural adaptation. Additional relevant papers from other sources are also discussed in this work to a lesser extent and coverage. Evaluation method and probing method papers are discussed to a much lesser extent due to the scope of our paper. We also refer readers to recent surveys like Adilazuarda et al. (2024) for focused discussions.